The point (2, 1) is a solution to which of the following systems of equations?
a. 2x + y = 5 and −2x + y = 2
b. −2x − y = −5 and 2x − y = 2
c. −x + y = 3 and −3x + y = −5
d. x + y = 3 and 3x − y = 5

Answer by ewatrrr(24785)


Hi
x + y = 3, 2 + 1 = 3
3x − y = 5, 6 - 1 = 5